# Criticality-Guided Efficient Pruning in Spiking Neural Networks Inspired by Critical Brain Hypothesis

### v2

@@ 7 Conclusion
In this paper, inspired by the critical brain hypothesis in neuroscience, we propose a novel neuron criticality metric and design a regeneration based on criticality for high-efficiency SNN pruning. Experimental results demonstrate that our method outperforms previous SNN pruning methods in both performance and efficiency for unstructured and structured pruning. Moreover, further experiments confirm our insight that the high criticality of the pruned model improves feature extraction. An interesting future direction is to further investigate the criticality in different scales of SNNs and exploit it to guide model design and high-efficiency learning.

### v3

@@ 7 Conclusion
In this paper, inspired by the critical brain hypothesis in neuroscience and the high biological plausibility of SNNs, we analyze criticality in SNNs from the perspective of maximizing feature information entropy and propose a low-cost metric for assessing neuron criticality. Incorporating this criticality into the pruning process facilitates the development of a highly efficient pruning-regeneration method. Experimental results demonstrate that our method outperforms previous SNN pruning methods in both performance and efficiency for unstructured and structured pruning. Moreover, further experiments confirm our insight that the high criticality of the pruned model improves feature extraction. An interesting future direction is to further explore the criticality in different scales of SNNs and exploit it to guide model design and high-efficiency learning.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (No. 41930110).